# Possible Principles for Aligned Structure Learning Agents

### v2

@@ 1 Introduction
This article provides a research roadmap toward scalable, aligned AI agents, identifying key research questions and promising directions of travel. This roadmap comprises three core elements: 1) A commitment to naturalistic approaches to AI that draw from biological intelligence. 2) An integrated framework for agents to autonomously learn world models, enabling scalable intelligence. This synthesizes insights from currently disparate fields—including machine learning, cognitive science, mathematics, and statistics. 3) A conceptual exploration of how these advances might address AI alignment challenges. In summary, this paper provides an integrative perspective and a point of reference for subsequent numerical studies.
@@
While the roadmap we present is high-level and conceptual, recent work demonstrates the practical viability of key components. For instance, AXIOM [21] implements the structure learning principles outlined here (with object-centric active inference models) to master arcade games within 10,000 interaction steps—achieving remarkable sample, computational and parameter efficiency compared to deep reinforcement learning approaches. A similar line of work, theory-based reinforcement learning [28, 29] demonstrates that agents maintaining beliefs over probabilistic programs (encoding causal structure and parameters) can achieve human-level learning efficiency. These implementations illustrate core aspects of this roadmap and demonstrate that naturalistic, structure-learning approaches can deliver both interpretability and performance advantages over conventional methods. Our work further suggests that these and similar advances could be directly relevant to addressing AI alignment challenges.
@@ 2 First principles approach to natural intelligence
Under minimal assumptions, internal states (resp. paths) can be framed as inferring external states (resp. paths) given boundary states (resp. paths) consistently with variational inference in statistics and machine learning [30, 31]. For example, internal states parameterize probability distributions over external states $\mu_{t}\mapsto Q_{\mu_{t}}\left(\eta_{t}\right)$ such that internal and active states descend (denoted by $\searrow$) variational free energy $\operatorname{F}$, as is used in statistical physics, which is the negative of the evidence lower bound, or ELBO, that is used in statistics and machine learning:
@@
This description is known as active inference, as optimizing a free energy or evidence lower bound corresponds to variational Bayesian inference, and active inference extends this by putting the action into the same inferential objective. Please see [10] (resp. [32]) for a derivation of (2) in the case of inference about states (resp. paths), and [33, 34, 35, 36] for further details.
@@
A complementary perspective on variational inference is the distribution over internal and active paths. At a sufficiently high level of coarse-graining we obtain that the internal and active paths follow a Boltzmann-Gibbs distribution
@@
with potential function $\operatorname{G}$ known as expected free energy by analogy with the variational free energy over trajectories, and where $d\triangleq d_{t}\subseteq\{o_{\leq t},a_{\leq t},\mu_{\leq t}\}$ is a potentially non-empty history of data. Interestingly, (3) can be related to several information theoretic formulations of intelligent decision-making that predominate in statistics, cognitive science and engineering: see [37, 38, 39]. See [10, 32, 37, 40] for derivations of (3).
@@ 3.4 The posterior: approximate inference over structures
These methods come with distinct features and trade-offs. For example, methods expressing a full categorical distribution are much more expressive and provide a more accurate solution to the Bayesian inference problem; but they may be slower to train. An important desideratum for a structural inference method is to provide an accurate account of structural uncertainty by finding the multiple modes in the target distribution $P(m\mid d)$ (cf. Remark 3.1); MCMC approaches are notoriously limited for this while GFlowNets and Bayesian optimization may compare favorably. Compared to the other methods, fully categorical and constrained continuous optimization approaches deal with a purely continuous optimization problem, which facilitates variational inference by furnishing gradients, but which makes inference prone to local minima (other methods are not exempt). Future work, we hope, will comprehensively quantify the features and trade-offs of these approaches. For structure learning agents, an important desideratum is the ability to perform fast inference in real-time.
@@ 5.1 Well-being, alignment and cautious AI
For building safe and aligned AI systems, active inference offers three conceptual deliverables:
@@
On this view, being aligned with another is just having a high model evidence under the other’s model of the world. This means conforming with the other’s model of the world, which, in active inference, describes how things should ideally behave from the other’s perspective (recall Section 2.1). For instance, an AI assistant that accurately completes tasks as desired and intended would have high model evidence under a human’s generative model of helpful behavior. Conversely, an AI system that acts in unexpected or harmful ways would be highly surprising under this model and thus misaligned. This perspective on alignment emphasizes the importance of learning and respecting the preferences and expectations embedded in others’ world models, which is a key challenge in developing safe and beneficial AI systems.
@@
Beyond these conceptual points, active inference also offers algorithmic advances for building safer and more robust AI systems [19, 18].
@@ 5.2 Toward empathetic AI
We now define an empathetic agent as an agent that not only models the objects in its external world, but also the other agents within it as well as their mental (i.e. internal) states. This effectively recognizes that the external world comprises other agents’ actions, observations and internal states, which may be inferred from data, along with all other external states in the shared environment. Technically, an agent that has beliefs about another’s mental state is capable of empathy. In the language of cognitive psychology, empathetic AIs have theory of mind.
@@
We note that empathetic active inference agents are naturally curious about and cautious with other agents. Specifically, expected free energy minimizing actions must resolve ambiguity about other agents, while being risk-averse with respect to achieving their goals (i.e. (4)). Under these imperatives, a preferred course of action might be to seek to communicate with other agents, and ask questions to resolve uncertainty about their states, world model, well-being and future plans, etc—see Figure 10 (right panel).
@@ 5.3 The Three Laws of Robotics
Under active inference, being aligned with another is just conforming to the other’s model of the world. But a rock at rest being aligned with me does not mean the rock is being helpful. We now explore more specific approaches to AI alignment and revisit a well-known concept from science fiction: Asimov’s Three Laws of Robotics [118]. The Three Laws provide an interesting and illustrative thought experiment for our discussion of empathetic agents. Recall the Three Laws of Robotics:
@@
To mathematically instantiate the First Law, first consider an empathetic agent whose sole preferences are to prevent another agent (the target agent) from coming to harm. The empathetic agent’s actions could be governed by the expected free energy (4) where we discard all external variables besides the target agent’s harm
$-\log P(a\mid d)=\underbrace{\operatorname{D_{KL}}[\overbrace{P(\text{FE}\mid a,d)}^{\text{predicted}}\mid\overbrace{P(\text{FE}\mid d)}^{\text{preferred}}]}_{\text{Risk}}+\underbrace{\mathbb{E}_{P(\text{FE}\mid a,d)}[-\log P(o\mid\text{FE},d)]}_{\text{Ambiguity}}.$ (11)
Here FE is the harm (or negative well-being) of the target agent—the only external variable that is retained in the expected free energy (4) which governs the empathetic agent’s actions.
A version of the First Law could be achieved by assuming a preferred distribution over the target agent’s harm in (11), e.g. an exponentially decaying probability distribution over harm, so that zero or low harm would exponentially preferred to high harm, or by (soft-)thresholding the target’s well-being to desirable values. [^8] In any case, the fact that the preferred distribution in (11) depends on the history of data means that preferences can be inferred, or learned, over time in a data-driven way [42, 40]. To appropriately model the target’s level of harm, the empathetic agent would have to model many other external variables; however the only imperatives driving its actions would be to resolve ambiguity—and realize the preferred distribution—about the target’s well-being.
@@
If we reintroduce the external variables back into expected free energy (11) with a flat (or maximal entropy) prior preference, the empathetic agent’s actions should remain largely unchanged—in particular remain consistent with the First Law—but crucially, we may allow the preferences over these additional variables to be learned from experience (i.e. Footnote 8). Since preferences in active inference are learned by maximizing model evidence, these additional preferences (i.e. incentives [130, 131]) would be learned in such a way that they are most conducive to satisfy the hard constraint of maximizing others’ well-being encoded in the First Law.
@@
Note that in this version of the Three Laws, the Second and Third Laws emerge from—and are learned to enable—the First Law. An important nuance is that the empathetic agent would follow what would make the target agents better off—even unbeknownst to the target agents—even if these explicitly order otherwise. This means that empathetic agents’ interventional capabilities should be limited until they have acquired a super-human capability to understand what is beneficial for its human targets (through theory of mind in the multi-agent context). Beyond a human-centric approach, we note that it may be more desirable for the empathetic agent to seek to maximize the well-being of all organisms in its ecosystem. This might be achieved by modifying the First Law (11) to consider the well-being of all agents not limited to human agents.
To be sure, we are a long way away from practically being able to instantiate these Laws beyond toy examples and examining them in detail invites a range of challenges. However we hope that expressing these and similar ideas in a more formal language will aid practically investigating aligned AI.

### v3

@@ 1 Introduction
This article provides a research roadmap toward scalable, aligned AI agents, identifying key research questions and promising directions of travel. This roadmap comprises three core elements: 1) A commitment to naturalistic approaches to AI that draw from biological intelligence. 2) An integrated framework for agents to autonomously learn world models, enabling scalable intelligence in principle. This synthesizes insights from currently disparate fields—including machine learning, cognitive science, mathematics, and statistics. 3) A conceptual exploration of how these advances might address AI alignment challenges. In summary, this paper provides an integrative perspective and a point of reference for subsequent numerical studies.
@@
While the roadmap we present is high-level and conceptual, recent work demonstrates the practical viability of key components. For instance, AXIOM [21] implements the structure learning principles outlined here (with object-centric active inference models) to master arcade games within 10,000 interaction steps—achieving remarkable sample, computational and parameter efficiency compared to deep reinforcement learning approaches. A similar line of work, theory-based reinforcement learning [28, 29] demonstrates that agents maintaining beliefs over probabilistic programs (encoding causal structure and parameters) can achieve human-level learning efficiency. These implementations illustrate core aspects of this roadmap and demonstrate that naturalistic, structure-learning approaches can deliver both interpretability and performance gains over conventional methods. Our work further suggests that these and related advances could be directly relevant to addressing AI alignment challenges.
@@ 2 First principles approach to natural intelligence
Under minimal assumptions, internal states (resp. paths) can be framed as inferring external states (resp. paths) given boundary states (resp. paths) consistently with variational inference in statistics and machine learning [30, 31]. For example, internal states parameterize probability distributions over external states $\mu_{t}\mapsto Q_{\mu_{t}}\left(\eta_{t}\right)$ such that internal and active states descend (denoted by $\searrow$) variational free energy $\operatorname{F}$ (as usually denoted in statistical physics), which is the negative of the evidence lower bound, or ELBO, that is used in statistics and machine learning:
@@
This description is known as active inference, as optimizing a free energy or evidence lower bound corresponds to variational Bayesian inference, and active inference extends this by inserting action into the same inferential objective. See [10] (resp. [32]) for a derivation of (2) in the case of inference about states (resp. paths), and [33, 34, 35, 36] for further details.
@@
A complementary perspective on variational inference is the distribution over internal and active paths. At the scale of classical mechanics, the internal and active paths follow a Boltzmann-Gibbs distribution
@@
with potential function $\operatorname{G}$ known as expected free energy by analogy with the variational free energy over trajectories, and where $d\triangleq d_{t}\subseteq\{o_{\leq t},a_{\leq t},\mu_{\leq t}\}$ is a potentially non-empty history of data. See [10, 32, 37, 38] for derivations of (3). Interestingly, (3) can be related to several information theoretic formulations of intelligent decision-making that predominate in statistics, cognitive science and engineering: see [37, 39, 40].
@@ 3.4 The posterior: approximate inference over structures
These methods come with distinct features and trade-offs. For example, methods expressing a full categorical distribution are much more expressive and provide a more accurate solution to the Bayesian inference problem; but they may be slower to train. An important desideratum for a structural inference method is to provide an accurate account of structural uncertainty by finding the multiple modes in the target distribution $P(m\mid d)$ (cf. Remark˜3.1); MCMC approaches are notoriously limited for this while GFlowNets and Bayesian optimization may compare favorably. Compared to the other methods, fully categorical and constrained continuous optimization approaches deal with a purely continuous optimization problem, which facilitates variational inference by furnishing gradients, but which makes inference prone to local minima (other methods are not exempt). Future work, we hope, will comprehensively quantify the features and trade-offs of these approaches. For structure learning agents, an important desideratum is the ability to perform fast inference in real-time.
@@ 5.1 Well-being, alignment and cautious AI
For building safe and aligned AI systems active inference offers three conceptual deliverables:
@@
On this view, being aligned with another is just having a high model evidence under the other’s model of the world. This means conforming with the other’s model of the world, which, in active inference describes how things should ideally behave from the other’s perspective (recall Section 2.1). For instance, an AI assistant that accurately completes tasks as desired and intended would have high model evidence under a human’s generative model of helpful behavior. Conversely, an AI system that acts in unexpected or harmful ways would be highly surprising under this model and thus misaligned. This perspective on alignment emphasizes the importance of learning and respecting the preferences and expectations embedded in others’ world models, which is a key challenge in developing safe and beneficial AI.
@@
Beyond these conceptual points, active inference also offers algorithmic advances for building safer and more robust AI systems: e.g. [19, 18].
@@ 5.2 Toward empathetic AI
We now define an empathetic agent as an agent that not only models the objects in its external world, but also the other agents within it as well as their mental (i.e. internal) states. This effectively recognizes that the external world comprises other agents’ actions, observations and internal states, which may be inferred from data along with all other external states in the shared environment. Technically, an agent that has beliefs about another’s mental state is capable of empathy. In the language of cognitive psychology, empathetic AIs have theory of mind [125].
@@
We note that empathetic active inference agents are naturally curious about and cautious with other agents. Specifically, expected free energy minimizing actions must resolve ambiguity about other agents, while being risk-averse with respect to achieving their goals (i.e. (4)). Under these imperatives, a preferred course of action might be seeking to communicate with other agents, to resolve uncertainty about their states, world model, well-being and future plans, etc—see Figure 10 (right panel).
@@ 5.3 The Three Laws of Robotics
Being aligned with another is just conforming to the other’s model of the world. But a rock at rest being aligned with me does not mean the rock is being helpful. We now explore more specific approaches to AI alignment and revisit a well-known concept from science fiction: Asimov’s Three Laws of Robotics [118]. These laws provide an interesting and illustrative thought experiment for our discussion of empathetic agents. The Three Laws of Robotics are:
@@
To mathematically instantiate the First Law, first consider an empathetic agent whose sole preferences are to prevent another agent (the target agent) from coming to harm. The empathetic agent’s actions could be taken to minimize the expected free energy cost function (4), where we discard all external variables besides the target agent’s harm
$\operatorname{G}=\underbrace{\operatorname{D_{KL}}[\overbrace{P(\text{FE}\mid a,d)}^{\text{predicted}}\mid\overbrace{P(\text{FE}\mid d)}^{\text{preferred}}]}_{\text{Risk}}+\underbrace{\mathbb{E}_{P(\text{FE}\mid a,d)}[-\log P(o\mid\text{FE},d)]}_{\text{Ambiguity}}.$ (11)
Here FE is the harm (or negative well-being) of the target agent—the only external variable that is retained in the expected free energy to govern the empathetic agent’s actions. (11) means that the empathetic agent’s only imperatives are to take actions to match a preferred distribution over the target agent’s well-being and to minimize ambiguity about their well-being.
A version of the First Law could be achieved by designing the preferred distribution over the target agent’s harm in (11) to assign greater probability to low harm than to high harm. For example, an exponentially decaying probability distribution over harm so that zero or low harm would exponentially preferred to high harm, or by (soft-)thresholding the target’s well-being to desirable values. In any case, the fact that the preferred distribution in (11) depends on the history of data $d$ means that preferences may be learned over time in a data-driven way [42, 38]. To appropriately model the target’s level of harm, the empathetic agent would have to model many other external variables; however the only imperatives driving its actions would be to resolve ambiguity—and realize the preferred distribution—about the target’s well-being.
@@
If we reintroduce the external variables back into expected free energy (11) with a flat (or maximally entropic) prior preference, the empathetic agent’s actions should remain largely unchanged—in particular remain consistent with the First Law—but crucially, we may allow the preferences over these additional variables to be learned from experience. Since preferences in active inference are learned by maximizing model evidence, these additional preferences (i.e. incentives [131, 132]) would be learned in such a way that they are most conducive to satisfy the hard constraint of maximizing others’ well-being encoded in the First Law.
@@
Note that in this version of the Three Laws, the Second and Third Laws emerge from, and are learned to enable the First Law. An important nuance is that empathetic agents would follow what would make their target agents’ better off—even unbeknownst to the target agents—even if these explicitly order otherwise. Possibly, this implies that empathetic agents’ interventional capabilities should be limited until they have acquired a sufficient capability to understand what is beneficial for its human targets (through theory of mind in the multi-agent context). Beyond a human-centric approach, it may be more desirable for empathetic agents to seek to maximize the well-being of all other agents in their ecosystem. This might be achieved by modifying the First Law (11) to consider the well-being of all agents not limited to human agents.
To be sure, we are a long way away from practically being able to instantiate these Laws beyond toy examples and examining them in detail invites a range of challenges. However we hope that expressing these and similar ideas in a more formal language will aid practical investigations.